OpenStudy (anonymous):

A circle has a radius of 18 inches. What is the circumference of the circle in terms of Pi Explain how you got your answer

OpenStudy (anonymous):

c=pi*diameter diameter=2*r are you ok now?

Parth (parthkohli):

36pi is the answer as circumference = 2*pi*r = 2*18*pi = 36pi